find the fifth term of the binomial expansion
(x^4+y^3)^8

Answer by nyc_function(2741)

The index k of each term
is one less than the ordinal number of the term.
Here is the formula you need:
(x + y)^n = (n over k) * x^(n - k) * y^k
Let x = x^4
Let y = y^3
Let k = 0, 1, 2, 3 and 4
Can you do it now?